Solid copper can be produced by passing gaseous ammonia over solid copper (II) oxide at high temperatures.

The other products of the reaction are nitrogen gas and water vapor. The balanced equation for this reaction is:

2NH3(g)+3CuO(s)->N2(g)+3Cu(s)+3H2O(g)

What is the theoretical yield of solid copper that should form when 18.1 g of NH3 is reacted with 90.4 g of CuO? If only 45.3 g of copper is actually collected, what is the percent yield?
Answer #1

Theoretical yield for

m = 18.1 g of NH3

mol = mass/MW = 18.1/17 = 1.0647 mol of NH3

m = 90.4 g of CuO

mol = mass/MW = 90.4 /79.545 = 1.1364 mol of CuO

ratio is

2:3 so

1.0647 mol of NH3 needs 3/2*1.0647= 1.59705 mol ofCUO, which we do not hav,e thereofre CuO is limiting

CuO: Cu ratio is 1:1 so

1.0647 mol of CuO will produce 1.0647 mol of Cu(s)

in mass:

mass of Cu(s) = mol*MW = 1.0647 *63.5 = 67.6084 g of Cu

b)

if only m = 45.3 g of Cu is actually collected then

% yield = real /theoreteical * 100 = 45.3 /67.6084 * 100 = 67.003508

% yield = 67.003508 %